State whether the following are true or false. Justify your answer. 
 
 The value of cos θ increases as θ increases.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The answer is False.

According to the values obtained as per the unit circle,
the values of cos are: cos 0° = 1

cos 30° = √3/2 cos 45° = 1/√2

cos 60° = 1

cos 90° = 0
Thus, the value ofcos θ decreases as θ increases.
So, the statement given above is false.

Thanks